how many molecules are in 2.7 moles?
1
Expert's answer
2020-10-08T13:55:18-0400

Solution:

The Avogadro constant NA= 6.022 × 1023 units in one mole of any substance. The units may be electrons, atoms, ions, or molecules.

The number of particles of a substance can be calculated using the Avogadro constant NA, the amount of substance in mole n.

Number of particles N = Avogadro constant NA × amount n

N = 6.022×1023 mol-1 × 2.7 mol = 16.259×1023 = 1.626×1024

Answer: there are 1.626×1024 molecules in 2.7 moles.